Give a numerical example of cost function and analyze this cost function? Discuss how is this cost function used in decision making?
Answer #1

Ans. Cost is the total cost of producing output

The cost function consists of two different type of cost

Variable cost

Fixed cost

Variable cost: variable cost varies with output. variable cost can be expressed as the product total variable cost per unit.

fixed cost: normally do not vary with output. in general this cost would be incurred whether the items are produced or not.

Cost function

C(x)   = F+Vx                   C= Total cost

                                  F= Fixed cost

                                  V= Variable cost

                                  x = no of unit produced

Example: Output of the produced 100units and Direct material $5 and direct labour is 2 and other direct variable exp. $1 and fixed cost during the period 10000.

Cost function

Calculation of total cost

Variable cost = 100X(5+2+1) = $800

Fixed cost     = 10000

calculation of total cost = 800+10000 = $10800

How to use cost function in decision making: With cost function we can analyze total cost of production and per unit cost of product and can also fix the selling price.

Cost function help the management for decision making for fixation selling price and indentifying cost area where improvement is required and reduced the waste or scrap.

Cost function help the presentation of cost related data in better way to help the management for decision making.